can you help me solve:
3-2(3-g) = 4(3+2g) -3g

Answer by nerdybill(7384)

3-2(3-g) = 4(3+2g) -3g
3-6+2g = 12+8g -3g
-3+2g = 12+5g
-3 = 12+3g
-15 = 3g
-5 = g